# Joyce structures from quadratic differentials on the sphere

### v1

@@ 4. Quadratic differentials and the moduli space $M={\rm Quad}(\bar{m})$
$=\sum_{a=1}^{n}\frac{\partial z_{a}}{\partial a_{i}^{(\alpha)}}\frac{\partial z_{a+n}}{\partial a_{j}^{(\beta)}}-\frac{\partial z_{a+n}}{\partial a_{i}^{(\alpha)}}\frac{\partial z_{a}}{\partial a_{j}^{(\beta)}}=\sum_{a=1}^{2n}\sum_{b=1}^{2n}\frac{\partial z_{a}}{\partial a_{i}^{(\alpha)}}\omega_{ab}\frac{\partial z_{b}}{\partial a_{j}^{(\beta)}}$ (4.27)
@@
The lattice $\mathcal{G}\to TM$ is given by the span of vector fields with integer coefficients with respect to the $\{z_{a}\}_{a=1}^{2n}$ coordinate basis. The coordinates $\{z_{a}\}_{a=1}^{2n}$ are therefore flat for the associated connection $\nabla$. The vector field $E_{0}$ defined by
$E_{0}=z_{a}\frac{\partial}{\partial z_{a}}$ (4.29)
@@ 5. Potentials and the complex manifold $X$
$R(x)=\frac{\displaystyle{\sum_{I=1}^{n}F(q_{I})\Big{(}2p_{I}v_{I}-\sum_{K\neq I}^{n}\frac{p_{K}}{q_{I}-q_{K}}\Big{)}\prod_{J\neq I}^{n}\frac{(x-q_{J})}{(q_{I}-q_{J})}}}{F(x)}$ (5.4)
@@
The remaining term in the potential (5.1) is completely determined by the point in $X$, if we impose a condition that the potential $Q(x)$ has a particular form of an apparent singularity at $x=q_{I}$. The condition to impose is that $Q(x)$ has Laurent expansions
@@
for some functions $u_{I}$ on $X$. That the existence of apparent singularities is necessary for the existence of isomonodromic deformations of scalar ODEs was shown by Poincaré [24]. The monodromy of the equation (1.1) is fixed at these points in such a way analytic continuation around the singularity changes the sign of the solution.
The apparent singularity condition forces
$Q_{2}(x)=\sum_{I=1}^{n}\frac{3}{4(x-q_{I})^{2}}+\sum_{I=1}^{n}\frac{v_{I}}{x-q_{I}}+S(x)$ (5.7)
@@
We consider the equation $y^{2}=Q(x)$ determined by a $\xi\in X$ and $\hbar\in\mathbb{C}^{*}$. In the Appendix B we prove a fact that we will need later, that for $(\xi,\hbar)\in X\times\mathbb{C}^{*}$ in the complement of the vanishing set of a non-zero holomorphic function $\Delta:X\times\mathbb{C}^{*}\to\mathbb{C}$, $Q(x)$ has simple zeroes. The Riemann-Hurwitz formula implies that, given $Q(x)$ has simple zeroes, the genus of the smooth algebraic curve $\Sigma(\xi,\hbar)$ defined by $y^{2}=Q(x)$ has genus $2n$.
@@
$X_{p}\cong(\mathcal{S}_{0}(p)\setminus\mathcal{B}(p)\times...\times\mathcal{S}_{0}(p)\setminus\mathcal{B}(p))\setminus\delta\times\mathbb{C}^{n}.$ (5.9)
@@ 6. Isomonodromic deformations and the cohomology of $\Sigma(\xi,\hbar)$
for some meromorphic functions $f_{I}(t)$.
@@
If the leading term of the Laurent expansion of $A$ at the point $x=q_{I}$ is $(x-q_{I})^{-M_{I}}$, $M_{I}\geq 1$ then equating coefficients of the $(x-q_{I})^{-M_{I}-3}$ term in the Laurent expansion of (6.1) yields
@@
The $x^{2m_{\infty}+M_{\infty}-6}$ term gives
$-4M_{\infty}-2(2m_{\infty}-5)=0.$ (6.10)
@@
$\Bigg{[}\frac{dy}{dt}dx\Bigg{]}=-[\kappa(A)]\in H^{1}(\Sigma(\xi,\hbar),\mathbb{C})$ (6.12)
@@
Similarly to §4 we use the notation $U(\Psi)]\coloneqq U(y)dx$. Consider now the linear map
@@
The right-hand side of the formula (6.15) is well-defined as a sum of residues at some poles $\{w_{\alpha},q_{I}\}$ of locally defined meromorphic functions for all $\xi\in X$ and $\hbar\in\mathbb{C}^{*}$. Therefore, for each $\hbar\in\mathbb{C}^{*}$ we may define a holomorphic $2$-form $\Omega$ on $X$ that agrees with the pullback of the intersection pairing $\langle\cdot,\cdot\rangle$ where $\Sigma(\xi,\hbar)$ is non-degenerate. In Appendix B we show that $\Sigma(\xi,\hbar)$ is non-degenerate on an open dense subset of $X\times\mathbb{C}^{*}$.
@@
The condition $L(\hbar)\subseteq\operatorname{ker}\Omega$ is a closed one so it suffices to show this for $(\xi,\hbar)$ for which the curve $\Sigma(\xi,\hbar)$ is non-singular. For such a $(\xi,\hbar)$ the intersection pairing $\langle\cdot|\cdot\rangle$ is of course non-degenerate. This implies $\operatorname{ker}\Omega=\mu(\xi,\hbar)^{-1}{((\operatorname{im}\mu(\xi,\hbar)})^{\perp})$. It therefore, using (6.12), suffices to show that $\kappa(A)]$ is orthogonal to $\operatorname{im}\mu(\xi,\hbar)$ with respect to $\langle\cdot|\cdot\rangle$ for all $A$ as in Lemma 6.1. Let $\kappa_{I}\coloneqq\kappa(A)$ for $A=(x-q_{I})^{-1}$. From the apparent singularity condition (5.6) at $x=q_{J}$ it follows
$y=\frac{\sqrt{3}}{2(x-q_{J})}+\frac{u_{J}}{\sqrt{3}}+\frac{2u_{J}^{2}}{3\sqrt{3}}(x-q_{J})+\frac{9P_{J}(q_{J})-4u_{J}^{3}}{9\sqrt{3}}(x-q_{J})^{2}+$... (6.16)
@@
$\kappa_{I}=\Bigg{(}-\frac{\sqrt{3}}{(x-q_{I})^{3}}+\frac{2u_{I}}{\sqrt{3}(x-q_{I})^{2}}+\frac{18P_{I}(q_{I})-16u_{I}^{3}}{9\sqrt{3}}+...\Bigg{)}dx.$ (6.17)
@@
We use these to calculate, for $U\in T_{\xi}X$ with no $\frac{\partial}{\partial q_{J}}$ terms for J=1,,nJ=1,...,n,
$d^{-1}\big{(}U(\Psi)\big{)}$ $=\frac{W(u_{J})}{\sqrt{3}}(x-q_{J})+\frac{2u_{J}U(u_{J})}{3\sqrt{3}}(x-q_{J})^{2}+O((x-q_{J})^{3}).$ (6.18)
@@
The above expansions imply for arbitrary $U\in T_{\xi}X$ that the $1$-form $\rho\coloneqq d^{-1}\big{(}U(\Psi)\big{)}\kappa_{I}$ has no residues at the pair of poles corresponding to $x=q_{I}$.
Lastly, we argue $\rho$ has no other poles. Since $k_{I}$ has a simple zero at $x=q_{J}$ for $J\neq I$ we see from above that these are not poles of $\rho$. That the poles of $\phi$ lying in $\mathbb{C}$ are not poles of $\rho$ follows because the function $d^{-1}(U(\Psi))$ has a pole of order at most $2m_{\alpha}-1$ at such a point while $\kappa_{I}$ has a zero of order $2m_{\alpha}$. At $x=\infty$, $d^{-1}(U(\Psi))$ has a pole of order at most $2m_{\infty}-7$ while $\kappa_{I}$ has a zero of order $2m_{\infty}$. ∎
@@ 7. The twistor distribution of isomonodromic flows
A subset of the isomonodromic flows will be “trivially isomonodromic” in the sense that they satisfy (6.1) with $A=0$. It is convenient to introduce some terminology:
@@
$L_{i}^{(\alpha)}\coloneqq\frac{\partial}{\partial a_{i}^{(\alpha)}}-\frac{1}{\hbar}\sum_{I=1}^{n}\frac{\partial p_{I}}{\partial a_{i}^{(\alpha)}}\frac{\partial}{\partial v_{I}}.$ (7.2)
@@
$m_{\infty}-4+\sum_{\alpha=1}^{N}m_{\alpha}=n-1$ (7.4)
@@
$L_{I}=U_{I}+\frac{V_{I}}{\hbar}$ (7.5)
with $U_{I},V_{I}$ given in Appendix A.
@@
Note the $(x-q_{I})^{0}$ and $(x-q_{I})^{-2}$ equations are implied by the other two. each For $K\neq I$, a more complicated computation gives
@@
The conditions give expressions for $\dot{q_{K}}$ and $\dot{u_{K}}$ for K=1,,nK=1,...,n. In particular, $\dot{v_{K}}$ is determined given $\dot{p_{K}}$. Altogether
@@
We have,
$P_{I}(x)=\frac{Q_{0}^{\prime}(x)}{\hbar^{2}}+\frac{R^{\prime}(x)-\sum\limits_{J\neq I}^{n}\frac{p_{j}}{(x-q_{j})^{2}}}{\hbar}-\sum_{J\neq I}^{n}\frac{3}{2(x-q_{j})^{3}}-\sum_{J\neq I}^{n}\frac{v_{J}}{(x-q_{J})^{2}}+S^{\prime}(x),$ (7.14)
while for K=1,,nK=1,...,n
@@
Substituting these into (7.12) and (7.13) we obtain expressions for $\dot{v_{K}}$ for K=1,,nK=1,...,n as functions of $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$ and $\{\dot{a_{i}}^{(\alpha)},\dot{w_{\alpha}}\}$.
@@
$-2(2m_{\alpha}-1)\frac{a_{2m_{\alpha}-1}^{(\alpha)}\dot{w_{\alpha}}}{(x-w_{\alpha})^{2m_{\alpha}}}=\frac{2(2m_{\alpha}-1)}{w_{\alpha}-q_{I}}\frac{a_{2m_{\alpha}-1}^{(\alpha)}}{(x-w_{\alpha})^{2m_{\alpha}}}$ (7.18)
@@
${\dot{a_{i}}^{(\alpha)}}={T^{(\alpha)}_{i}(q_{I}-w_{\alpha})}$ (7.20)
@@
$T_{i}^{(\alpha)}(x)=\sum_{k=i}^{2m_{\alpha}-1}(k-2)(k-i+1)!\frac{a_{k}^{(\alpha)}}{x^{k-i+2}}.$ (7.21)
@@
We have now specified $\{\dot{a_{i}}^{(\alpha)},\dot{w_{\alpha}},\dot{v_{I}},\dot{q_{I}}\}$ in terms of $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$ in such a way that that (6.1) is satisfied up to a rational function with roots at the $n$ points $q_{i}$ but with poles the orders of which sum to at most
@@
Thus this rational function is zero and (6.1) is satisfied. We therefore see that (given the freedom to add isopotential flows) there is an $n+1$ dimensional space of flows satisfying (6.1) with $A=(x-q_{I})^{-1}$ for each $I$. We may of course add the unique isopotential flow so that $\dot{a_{i}}^{(\alpha)}=0$ where $1\leq i\leq m_{\alpha}$ for $\alpha=1,...,N$ and $0\leq i\leq m_{\infty}-4$ for $\alpha=\infty$. We denote the generator of this distinguished flow, with $A$ as above, by $L_{I}$.
Note that, substituting everything into (7.12) and (7.13), the expressions for the time derivatives $\{\dot{a_{i}}^{(\alpha)},\dot{w_{\alpha}},\dot{v_{I}},\dot{q_{I}}\}$ in terms of $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$ contain only $\hbar$ and $\hbar^{-1}$ terms. In particular, we may write
@@
The full set of isomonodromic flows $\{L_{i}^{(\alpha)}\}\cup\{L_{I}\}$ are given explicitly in Appendix A. It remains to show that the set of $4n$ vector fields
@@
trivialise TXTX and so define a twistor distribution in the sense of Definition 2.2. This is essentially a combinatorial argument and so deferred to Appendix C. ∎

### v2

@@ 4. Quadratic differentials and the moduli space $M={\rm Quad}(\bar{m})$
$=\sum_{a=1}^{n}\frac{\partial z^{a}}{\partial a_{i}^{(\alpha)}}\frac{\partial z^{a+n}}{\partial a_{j}^{(\beta)}}-\frac{\partial z^{a+n}}{\partial a_{i}^{(\alpha)}}\frac{\partial z^{a}}{\partial a_{j}^{(\beta)}}=\sum_{a=1}^{2n}\sum_{b=1}^{2n}\frac{\partial z^{a}}{\partial a_{i}^{(\alpha)}}\omega_{ab}\frac{\partial z^{b}}{\partial a_{j}^{(\beta)}},$ (4.27)
@@
The lattice $\mathcal{G}\to TM$ is given by the span of vector fields with integer coefficients with respect to the $\{z^{a}\}_{a=1}^{2n}$ coordinate basis. The coordinates $\{z^{a}\}_{a=1}^{2n}$ are therefore flat for the associated connection $\nabla$. The vector field $E_{0}$ defined by
$E_{0}=\sum_{a=1}^{2n}z^{a}\frac{\partial}{\partial z^{a}}$ (4.29)
@@
Remark 4.6.
The above construction differs from the conjectured general construction of Joyce structures of class $S[A_{1}]$ outlined in [10], which uses the natural period structure on $M$ coming from Donaldson-Thomas theory. The difference comes from considering integral homology of the curve $\Sigma_{0}(p)$ punctured at the set of poles $\mathcal{P}_{0}(p)$ of $\psi$. Our lattice is dual to the integral cycles $\{\gamma_{a}\}$ defining a basis for $H_{1}(\Sigma_{0}(p),\mathbb{Z})$ whereas the integral lattice in the proposed construction is dual to a basis for $H_{1}^{-}(\Sigma_{0}(p)\setminus\mathcal{P}_{0}(p),\mathbb{Z})$, which is the subgroup of $H_{1}(\Sigma_{0}(p)\setminus\mathcal{P}_{0}(p),\mathbb{Z})$ consisting of classes anti-invariant under the covering involution. The natural homomorphism $H_{1}^{-}(\Sigma_{0}(p)\setminus\mathcal{P}_{0}(p),\mathbb{Z})\to H_{1}(\Sigma_{0}(p),\mathbb{Z})$ induced by the inclusion $\Sigma_{0}(p)\setminus\mathcal{P}_{0}(p)\hookrightarrow\Sigma_{0}(p)$ is injective and of full-rank but not in general surjective.
The resulting Joyce structures will share the skew-form $\omega$ on $M$, underlying hyper-Kähler metric and Euler vector field, but the period coordinates will be different, related by a constant linear transformation.
@@ 5. Potentials and the complex manifold $X$
$R(x)=\frac{\displaystyle{\sum_{I=1}^{n}F(q_{I})\Big(2p_{I}v_{I}-\sum_{K\neq I}^{n}\frac{p_{K}}{q_{I}-q_{K}}\Big)\prod_{J\neq I}^{n}\frac{(x-q_{J})}{(q_{I}-q_{J})}}}{F(x)},$ (5.4)
@@
The remaining term in the potential (5.1) is completely determined by the point in $X$ if we impose two conditions on $Q(x)$. The first is that $Q(x)$ has a particular form of an apparent singularity at $x=q_{I}$. It is that $Q(x)$ has Laurent expansions
@@
for some functions $u_{I}$ on $X$. The second condition is that $Q_{2}(x)$, as with $Q_{1}(x)$, has poles at the poles of $Q_{0}(x)$ in such a way that $Q_{2}(x)/2y_{0}\,dx$ has poles only at the points corresponding to $x=q_{I}$.
That the existence of apparent singularities is necessary for the existence of isomonodromic deformations of scalar ODEs was shown by Poincaré [27]. The monodromy of the equation (1.1) is fixed at these points in such a way that analytic continuation around the singularity changes the sign of the solution.
The reason for the choice of parameters $\{v_{I}\}$ is that
$Q_{1}(x):=\frac{p_{I}}{x-q_{I}}+2p_{I}v_{I}+O(x-q_{I}).$ (5.7)
Then, since $Q_{0}(x)=p_{I}^{2}+O(x-q_{I})$, the condition (5.6) implies $u_{I}=p_{I}/\hbar+v_{I}$ and the two conditions together force
$Q_{2}(x)=\sum_{I=1}^{n}\frac{3}{4(x-q_{I})^{2}}+\sum_{I=1}^{n}\frac{v_{I}}{x-q_{I}}+S(x),$ (5.8)
@@
We consider the equation $y^{2}=Q(x)$ determined by a $\xi\in X$ and $\hbar\in\mathbb{C}^{*}$. In Appendix C, we prove a fact that we will need later, that for $(\xi,\hbar)\in X\times\mathbb{C}^{*}$ in the complement of the vanishing set of a non-zero holomorphic function $\Delta:X\times\mathbb{C}^{*}\to\mathbb{C}$, $Q(x)$ has simple zeroes. The Riemann-Hurwitz formula implies that, given $Q(x)$ has simple zeroes, the genus of the smooth algebraic curve $\Sigma(\xi,\hbar)$ defined by $y^{2}=Q(x)$ is $2n$.
@@
$X_{p}\cong(\mathcal{S}_{0}(p)\setminus\mathcal{B}(p)\times...\times\mathcal{S}_{0}(p)\setminus\mathcal{B}(p))\setminus\delta\times\mathbb{C}^{n}$ (5.10)
@@ 6. Isomonodromic deformations and the cohomology of $\Sigma(\xi,\hbar)$
for some holomorphic functions $f_{I}(t)$.
@@
If the leading term of the Laurent expansion of $A$ at the point $x=q_{I}$ is, up to scaling by a constant, $(x-q_{I})^{-M_{I}}$, $M_{I}\geq 1$, then equating coefficients of the $(x-q_{I})^{-M_{I}-3}$ term in the Laurent expansion of (6.1) yields
@@
Recall $m_{\infty}\geq 3$. The $x^{2m_{\infty}+M_{\infty}-6}$ term gives
$-4M_{\infty}-2(2m_{\infty}-5)=0$ (6.10)
@@
$\Bigg[\frac{dy}{dt}dx\Bigg]=-[\kappa(A)]\in H^{1}(\Sigma(\xi,\hbar),\mathbb{C}),$ (6.12)
@@
Similarly to §4 we use the notation $U(\Psi)\coloneqq U(y)dx$. Consider now the linear map
@@
The right-hand side of the formula (6.15) is well-defined as a sum of residues at some poles $\{w_{\alpha},q_{I}\}$ of locally defined meromorphic functions for all $\xi\in X$ and $\hbar\in\mathbb{C}^{*}$. Therefore, for each $\hbar\in\mathbb{C}^{*}$ we may define a holomorphic $2$-form $\Omega$ on $X$ that agrees with the pullback of the intersection pairing $\langle\cdot|\cdot\rangle$ where $\Sigma(\xi,\hbar)$ is non-degenerate. In Appendix C we show that $\Sigma(\xi,\hbar)$ is non-degenerate on an open dense subset of $X\times\mathbb{C}^{*}$.
@@
The condition $L(\hbar)\subseteq\operatorname{ker}\Omega$ is a closed one so it suffices to show this for $(\xi,\hbar)$ for which the curve $\Sigma(\xi,\hbar)$ is non-singular. For such a $(\xi,\hbar)$ the intersection pairing $\langle\cdot|\cdot\rangle$ is of course non-degenerate. This implies $\operatorname{ker}\Omega=\mu(\xi,\hbar)^{-1}{((\operatorname{im}\mu(\xi,\hbar)})^{\perp})$. We know from (6.12) that the image of the generator of an isomonodromic flow $V$ is the cohomology class $\kappa(A)]$ for some $A$ as in Lemma 6.1. We must therefore show that $\kappa(A)]$ is orthogonal to the image of $\mu(\xi,\hbar)$. By linearity, we may check this separately for each $\kappa_{I}\coloneqq\kappa(A)$ where $A=(x-q_{I})^{-1}$, for I=1,,nI=1,...,n.
From the apparent singularity condition (5.6) at $x=q_{J}$ it follows
$y=\frac{\sqrt{3}}{2(x-q_{J})}+\frac{u_{J}}{\sqrt{3}}+\frac{2u_{J}^{2}}{3\sqrt{3}}(x-q_{J})+\frac{9P_{J}(q_{J})-4u_{J}^{3}}{9\sqrt{3}}(x-q_{J})^{2}+$..., (6.16)
@@
$\kappa_{I}=\Bigg(-\frac{\sqrt{3}}{(x-q_{I})^{3}}+\frac{2u_{I}}{\sqrt{3}(x-q_{I})^{2}}+\frac{18P_{I}(q_{I})-16u_{I}^{3}}{9\sqrt{3}}+...\Bigg)dx,$ (6.17)
@@
We will now, for arbitrary $U\in T_{\xi}X$, show that the $1$-form $\rho_{I}\coloneqq d^{-1}U(\Psi)\kappa_{I}$ has no residues and hence that $\langle U(\Psi),\kappa_{I}\rangle=0$ as required.
We use (6.16) to calculate, for $U\in T_{\xi}X$ with no $\frac{\partial}{\partial q_{J}}$ terms for J=1,,nJ=1,...,n,
$d^{-1}\big(U(\Psi)\big)$ $=\frac{U(u_{J})}{\sqrt{3}}(x-q_{J})+\frac{2u_{J}U(u_{J})}{3\sqrt{3}}(x-q_{J})^{2}+O((x-q_{J})^{3}).$ (6.18)
@@
The above expansions imply for arbitrary $U\in T_{\xi}X$ that the $1$-form $\rho_{I}$ has no residues at the pair of poles corresponding to $x=q_{I}$.
Lastly, we argue $\rho_{I}$ has no other poles. Since $k_{I}$ has a simple zero at $x=q_{J}$ for $J\neq I$, we see from the above formulae that these are not poles of $\rho_{I}$. That the poles of $\phi$ lying in $\mathbb{C}$ are not poles of $\rho_{I}$ follows because the function $d^{-1}(U(\Psi))$ has a pole of order at most $2m_{\alpha}-1$ at such a point while $\kappa_{I}$ has a zero of order $2m_{\alpha}$. At $x=\infty$, $d^{-1}(U(\Psi))$ has a pole of order at most $2m_{\infty}-7$ while $\kappa_{I}$ has a zero of order $2m_{\infty}$. ∎
@@ 7. The twistor distribution of isomonodromic flows
A subset of the isomonodromic flows will be “trivially” isomonodromic in the sense that they satisfy (6.1) with $A=0$. It is convenient to introduce some terminology:
@@
$L_{i}^{(\alpha)}\coloneqq\frac{\partial}{\partial a_{i}^{(\alpha)}}-\frac{1}{\hbar}\sum_{I=1}^{n}\frac{\partial p_{I}}{\partial a_{i}^{(\alpha)}}\frac{\partial}{\partial v_{I}},$ (7.2)
where $1\leq i\leq m_{\alpha}$ and $\alpha=1,...,N$ and $0\leq i\leq m_{\infty}-4$ for $\alpha=\infty$.
@@
$m_{\infty}-4+\sum_{\alpha=1}^{N}m_{\alpha}=n-1,$ (7.4)
@@
$L_{I}=U_{I}+\frac{V_{I}}{\hbar},$ (7.5)
with $U_{I},V_{I}$ given in Appendix B.
@@
Note the $(x-q_{I})^{0}$ and $(x-q_{I})^{-2}$ equations are implied by the other two. For $K\neq I$, a more complicated computation gives
@@
The conditions give expressions for $\dot{q_{K}}$ and $\dot{u_{K}}$ for K=1,,nK=1,...,n. Altogether
@@
For K=1,,nK=1,...,n, a simple application of the chain rule yields:
@@
Meanwhile, in full detail:
$P_{I}(x)=\frac{Q_{0}^{\prime}(x)}{\hbar^{2}}+\frac{R^{\prime}(x)-\sum\limits_{J\neq I}^{n}\frac{p_{J}}{(x-q_{J})^{2}}}{\hbar}-\sum_{J\neq I}^{n}\frac{3}{2(x-q_{J})^{3}}-\sum_{J\neq I}^{n}\frac{v_{J}}{(x-q_{J})^{2}}+S^{\prime}(x).$ (7.17)
Substituting (7.15) through (7.17) into (7.12) and (7.13) we obtain expressions for $\dot{v_{K}}$ for K=1,,nK=1,...,n as functions of $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$ and $\{\dot{a_{i}}^{(\alpha)},\dot{w_{\alpha}}\}$.
@@
$-2(2m_{\alpha}-1){a_{2m_{\alpha}-1}^{(\alpha)}\dot{w_{\alpha}}}=\frac{2(2m_{\alpha}-1)}{w_{\alpha}-q_{I}}{a_{2m_{\alpha}-1}^{(\alpha)}}$ (7.18)
@@
${\dot{a_{i}}^{(\alpha)}}={T^{(\alpha)}_{i}(q_{I}-w_{\alpha})},$ (7.20)
@@
$T_{i}^{(\alpha)}(x)=\sum_{k=i}^{2m_{\alpha}-1}(2i-k-2)\frac{a_{k}^{(\alpha)}}{x^{k-i+2}}.$ (7.21)
@@
We have now specified $\{\dot{a_{i}}^{(\alpha)},\dot{w_{\alpha}},\dot{v_{I}},\dot{q_{I}}\}$ in terms of $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$ in such a way that (6.1) is satisfied up to a rational function with roots at the $n$ points $q_{I}$ but with poles the orders of which sum to at most
@@
Thus, this rational function is zero and (6.1) is satisfied. We therefore see that (given the freedom to add isopotential flows) there is an $n+1$ dimensional space of flows satisfying (6.1) with $A=(x-q_{I})^{-1}$ for each $I$. We may of course add the unique isopotential flow so that $\dot{a_{i}}^{(\alpha)}=0$ where $1\leq i\leq m_{\alpha}$ for $\alpha=1,...,N$ and $0\leq i\leq m_{\infty}-4$ for $\alpha=\infty$. We denote the generator of this distinguished flow, with $A$ as above, by $L_{I}$.
Note that, substituting everything into (7.12) and (7.13), the expressions for the time derivatives $\{\dot{a_{i}}^{(\alpha)},\dot{w_{\alpha}},\dot{v_{I}},\dot{q_{I}}\}$ in terms of $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$ contain only $\hbar^{0}$ and $\hbar^{-1}$ terms. In particular, we may write
@@
The full set of isomonodromic flows $\{L_{i}^{(\alpha)}\}\cup\{L_{I}\}$ are given explicitly in Appendix B. It remains to show that the set of $4n$ vector fields
@@
trivialise TXTX and so define a twistor distribution in the sense of Definition 2.2. This is essentially a combinatorial argument and so deferred to Appendix D. ∎